A basketball player scored 17 times during one game. He scored a total of 26 points, two for each two-point shot and one for each free throw. How many two-point shots dis her make? How many free throws?
Answer by TimothyLamb(4379)

t + f = 17
t = 17 - f
(2)t + (1)f = 26
---
(2)t + (1)f = 26
2t + f = 26
34 - 2f + f = 26
---
f = 8
t = 9
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php